# Observational imprints of tidal internal gravity wave dissipation in star-planet systems

### v1

@@ Abstract
Tidal interactions play a crucial role in the orbital evolution of close-in star-planet systems. There are numerous manifestations of tides, including planetary orbital migration, breaking resonant chains, tidal heating, orbital circularization, spin-orbit alignment, and stellar and planetary spin synchronization. In the present study, we focus on the dissipation of internal gravity waves within stars. We examine two mechanisms: wave breaking in stars with radiative cores and magnetic wave conversion in stars with convective cores. Applying tidal prescriptions modelling these processes, we demonstrate that the enhanced stellar rotation of both TOI-2458 and GJ 504 can be explained by the previous engulfment of a hot Jupiter caused by gravity wave damping. Furthermore, we show that the observed population of hot Jupiters can be divided into two distinct subsamples: those that are too young for gravity wave dissipation and those where it is ongoing. These subsamples exhibit qualitatively different orbital period distributions: young systems have a uniform distribution, while older systems show a steep decline at short orbital periods. Using a population synthesis approach, we successfully reproduce the main features of the older hot Jupiter sample based on the distribution of the younger systems. According to our estimates, up to 20% of the main-sequence stars within the mass range [0.7,1.5] $M_{\odot}$ that once hosted a hot Jupiter may have since engulfed it. Our results highlight the key role of internal gravity wave dissipation in shaping the orbital architectures of hot Jupiter systems.
@@ 2.1 Dissipation mechanisms
While equation (2) in principle applies if the waves are fully damped according to any mechanism (such as radiative diffusion of low frequency waves), we will apply it only when either the waves have large enough amplitudes to become strongly nonlinear and break, or via the linear process of magnetic wave conversion. We do not consider additional linear and weakly non-linear processes explicitly, such as those involving weaker radiative damping of IGWs (the neglect of which is more appropriate here than in more massive stars with radiative envelopes, e.g. Zahn, 1975, 1977; Savonije and Papaloizou, 1983; Goldreich and Nicholson, 1989) or global g-modes (Goodman and Dickson, 1998), and we also ignore excitation of secondary waves by weakly nonlinear interactions (Essick and Weinberg, 2016; Weinberg et al., 2024), which have been explored in some works in the literature. Studying their impact on star-planet orbital evolution is left to future work.
@@ 2.2 Orbital evolution
The parameter $f_{\mathrm{p}}$ depends on the planetary structure and constitution. In the present study, $f_{\mathrm{p}}$ is fixed at 2.4, which is approximately valid for a homogeneous fluid body. This value is higher than the value $f_{\mathrm{p}}=2.16$ in e.g. Faber et al. (2005); Ford and Rasio (2006); Nelson et al. (2017), but lower than the value $f_{\mathrm{p}}=2.7$ inferred from the three-dimensional hydrodynamical simulations of Guillochon et al. (2011).
@@ 2.3 Stellar models
Figure 1: Evolution of the critical magnetic field strength, $B_{\mathrm{crit}}$, for a 1.2 $M_{\odot}$ solar-metallicity model and tidal period of 0.5 days. Each color represents a different grid resolution, defined by the parameter mesh_delta_coeff, given in the legend. Solid (dotted) lines correspond to models without (with) element diffusion. Gray dashed line shows the radial magnetic field strength of the convective core, $B_{r}$.
In this study, we compute our stellar models using the evolution code MESA r11701 (Paxton et al., 2011, 2013, 2015, 2018, 2019) and inlist files from Gossage et al. (2021) based on the MIST framework (Dotter, 2016; Choi et al., 2016). We find that the critical magnetic field strength required for wave conversion, $B_{\mathrm{crit}}$, is sensitive to the grid resolution, which is determined by the parameter mesh_delta_coeff. Increasing the resolution (by decreasing mesh_delta_coeff) allows $B_{\mathrm{crit}}$ to converge at mesh_delta_coeff $\lesssim 0.5$, as we show in Fig.1. Therefore, we select mesh_delta_coeff =0.5=0.5 to balance computational efficiency and accuracy.
@@ 3.1 TOI-2458
Figure 3: Evolution of stellar rotation (top panel) and planetary orbital (bottom panel) periods for two star-planet systems, shown in blue and red. Left panel: $M_{\mathrm{pl}}=$ 0.7 and 4 $M_{\mathrm{J}}$, and $P_{\mathrm{orb}}=$ 1.5 and 2.5 days, respectively. Right panel: $M_{\mathrm{pl}}=$ 10 and 2 $M_{\mathrm{J}}$, and $P_{\mathrm{orb}}=$ 1 and 2.5 days, respectively. The black dotted line depicts the rotation period evolution of a solitary star. Circles and crosses denote the onset of IGW breaking and Roche-lobe overflow (and planetary engulfment), respectively. The green square with error bars shows TOI-2458’s observed rotation period and age, along with their uncertainties from Šubjak et al. (2025).
Our hypothesis that IGW dissipation can explain the star’s anomalously fast rotation is supported by Fig. 3. Its left panel illustrates the orbital evolution of two hypothetical HJs around TOI-2458, with planetary masses of 0.7 and 4 Jovian masses, and initial orbital periods of 1.5 and 2.5 days, respectively. The right panel illustrates the cases of systems with planetary masses of 10 and 2 Jovian masses, and initial orbital periods of 1 and 2.5 days, respectively. These initial orbital periods represent the range of possible separations between the engulfed planet and its host star, considering the proximity of the hot Neptune TOI-2458 b. For the least massive HJs, such as those with masses as small as $0.3M_{J}$, the smaller initial separations above would place them below the Roche limit (for $M_{\mathrm{pl}}=0.3\>M_{\mathrm{J}}$ and $R_{\mathrm{pl}}=1.5\>R_{\mathrm{J}}$, equation 9 yields $a_{\mathrm{R}}\approx 0.026$ AU, which corresponds to an orbital period of $\approx 1.5$ days) Conversely, a planet with a larger initial separation would be too close to TOI-2458 b, as the mutual separation between the two planets would fall below 4 mutual Hill radii, which is the stability threshold for two-planet systems proposed by Wu et al. (2019). The mutual Hill radius $\Delta_{\mathrm{H}}$ in the units of mutual planetary separation is determined by:
@@
Rapid orbital migration, triggered by IGW breaking (the onset of which is shown by the black circles), modifies the stellar rotation period compared to that of a solitary star (shown with a black dotted line), as depicted in the top panel of Fig. 3. The bottom left panel reveals that the engulfment of a 0.7 $M_{\mathrm{J}}$ planet occurs almost instantaneously after the onset of wave breaking. In contrast, a 4.0 $M_{\mathrm{J}}$ planet undergoes a prolonged phase of orbital migration that lasts more than 700 Myr. Despite the more massive planet depositing more angular momentum into stellar rotation, its orbital decay completes earlier, leaving more time for magnetic braking to erase the tidally enhanced rotation. Eventually, the rotation periods in both simulations converge, and at TOI-2458’s age, they reach values consistent with observations. Both of the additional example systems shown in the right panel also allow the star to match the rotation period measurements. However, this time, the stellar rotation period equals its observed value at ages corresponding to either the lower limit from Šubjak et al. (2025) or an upper limit based on the TAMS age (which exceeds the upper age limit from Šubjak et al. (2025); given that other parameters such as $T_{\mathrm{eff}}$ and log g suggest that TOI-2458 is an MS star, here we use TAMS as the upper age limit).
@@ 3.2 GJ 504
TOI-2458 and GJ 504 represent two possible examples where efficient IGW dissipation leading to the HJ engulfment can explain the faster rotation rates of their host stars over expectations for single stars. They thus provide possible evidence for both the wave breaking scenario in a solar-type star (for TOI-2458) and for the magnetic wave conversion scenario in a hotter star with a convective core (GJ 504).
@@ 4.5 Results
According to our synthetic populations, 18 – 20% of stars that once hosted a HJ and have not yet evolved off the MS have engulfed their HJ. The above fraction is calculated relative to all synthetic systems, including young and non-transiting ones. This finding is in good agreement with Lazovik (2023), although the latter study focused on solar-mass stars and considered a uniform SFH between 0 and 7 Gyr. Furthermore, we can estimate the relative number of HJs producing transit timing variations detectable with modern facilities over 10 years of observations. The cumulative shift in transit times $T_{\mathrm{shift}}$ can be derived from the following (e.g. Wilkins et al., 2017):
@@ 5.3 Other caveats
Another source of uncertainty is the contribution of planets that arrive at their present-day locations via high-eccentricity migration at late ages ($t>t_{\mathrm{gw}}$), which we refer to as ‘late-travelers’. Because our approach is designed to show a connection between the populations of young and old HJs, the late-stage implantation of planets would decrease the convergence between our synthetic and observed populations. Nevertheless, neglecting the contribution of these late-travelers is a reasonable simplification, as the dynamical instabilities that drive high-eccentricity migration are expected to take place relatively early in a system’s evolution. This is supported by observations of hot Neptunes, which were found to be more abundant in systems with ages between 100 Myr and 1 Gyr (Fernandes et al., 2025). It is possible, though, that a fraction of HJs are formed through secular interactions between planets on longer timescales, as discussed in e.g. Hamer and Schlaufman (2022).
@@ 6 Conclusions
  • Up to 20% of stars that once hosted a HJ and have not yet evolved off the MS have engulfed their planet.
@@ Acknowledgments
We gratefully acknowledge Dr. Chowbay Swastik for the provided data on ages and other parameters of star-planet systems. YL thanks Anastasiya Yarovova for helpful discussions. This research has used the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.

### v2

@@ Abstract
Tidal interactions play a crucial role in the orbital evolution of close-in star-planet systems. There are numerous manifestations of tides, including planetary orbital migration, breaking resonant chains, tidal heating, orbital circularization, spin-orbit alignment, and stellar and planetary spin synchronization. In the present study, we focus on the dissipation of internal gravity waves within stars. We examine two mechanisms: wave breaking in stars with radiative cores and magnetic wave conversion in stars with convective cores. Applying tidal prescriptions modelling these processes, we demonstrate that the enhanced stellar rotation of both TOI-2458 and GJ 504 can be explained by the previous engulfment of a hot Jupiter caused by gravity wave damping. Furthermore, we show that the observed population of hot Jupiters can be divided into two distinct subsamples: those that are too young for gravity wave dissipation and those where it is ongoing. These subsamples exhibit qualitatively different orbital period distributions: young systems have a uniform distribution, while older systems show a steep decline at short orbital periods. Using a population synthesis approach, we successfully reproduce the main features of the older hot Jupiter sample based on the distribution of the younger systems. According to our estimates, up to 13% of the main-sequence stars within the mass range [0.7,1.5] $M_{\odot}$ that once hosted a hot Jupiter may have since engulfed it. Our results highlight the key role of internal gravity wave dissipation in shaping the orbital architectures of hot Jupiter systems.
@@ 2.1 Dissipation mechanisms
While equation (2) in principle applies if the waves are fully damped according to any mechanism (such as radiative diffusion of low frequency waves), we will apply it only when either the waves have large enough amplitudes to become strongly nonlinear and break, or via the linear process of magnetic wave conversion. We do not consider additional linear and weakly non-linear processes explicitly, such as those involving weaker radiative damping of IGWs (the neglect of which is more appropriate here than in more massive stars with radiative envelopes, e.g. Zahn, 1975, 1977; Savonije and Papaloizou, 1983; Goldreich and Nicholson, 1989) or global g-modes (Goodman and Dickson, 1998), and we also ignore excitation of secondary waves by weakly nonlinear interactions (Essick and Weinberg, 2016; Weinberg et al., 2024), which have been explored in some works in the literature. Studying their impact on star-planet orbital evolution is left to future work. Furthermore, equation (2) assumes that the squared buoyancy frequency $N^{2}$ at the radiative/convective boundary depends linearly upon the distance from the boundary. As demonstrated by Barker (2011), accounting for deviations from this linear profile may introduce a factor of $\sim 2$ variation in $Q^{\prime}$. However, such differences would be insufficient to alter the primary results of this paper.
@@ 2.2 Orbital evolution
The parameter $f_{\mathrm{p}}$ depends on the planetary structure and constitution. In the present study, $f_{\mathrm{p}}$ is fixed at 2.4, which is approximately valid for a homogeneous fluid body. This value is higher than the value $f_{\mathrm{p}}=2.16$ in e.g. Faber et al. (2005); Ford and Rasio (2006); Nelson et al. (2017), but lower than the value $f_{\mathrm{p}}=2.7$ inferred from the three-dimensional hydrodynamical simulations of Guillochon et al. (2011). Although this work assumes tidal disruption upon filling the Roche lobe, another possibility is the stable accretion scenario proposed by Valsecchi et al. (2014) and further investigated by Lazovik (2023). According to this scenario, HJs can avoid immediate tidal disruption by continuously adjusting their radii to fill their Roche lobes while gradually losing mass, eventually transforming into hot Neptunes with thin envelopes, or Super-Earths. However, the validity of this scenario relies on an efficient mechanism to return the angular momentum of the transferred mass back to the planet’s orbit. The existence of such a mechanism was questioned by Jia and Spruit (2017), who argued that the star acts as a sink for the angular momentum, preventing orbital expansion and instead accelerating runaway mass transfer, which in the end causes tidal disruption. Recently, Hallatt and Millholland (2026) proposed an impulsive mass-loss regime, according to which a small return of angular momentum ($\sim 5-10\%$) is able to prevent engulfment, leading to the rapid shedding of most of the planet’s envelope. Nevertheless, even if a planet can survive the mass loss following Roche lobe overflow, we expect IGW dissipation to be sufficiently strong to subsequently drive inspiral into the star within a short timescale.
@@ 2.3 Stellar models
Figure 1: Evolution of the critical magnetic field strength, $B_{\mathrm{crit}}$, for a 1.2 $M_{\odot}$ solar-metallicity model and tidal period of 0.5 days. Each color represents a different grid resolution, defined by the parameter mesh_delta_coeff, given in the legend. $B_{\mathrm{crit}}$ is evaluated as a local minimum within the region $R_{\mathrm{c}},R_{\mathrm{c}}+0.2R_{*}$]
. Solid (dotted) lines correspond to models without (with) element diffusion. Gray dashed line shows the radial magnetic field strength of the convective core, $B_{r}$.
In this study, we compute our stellar models using the evolution code MESA r11701 (Paxton et al., 2011, 2013, 2015, 2018, 2019) and inlist files from Gossage et al. (2021) based on the MIST framework (Dotter, 2016; Choi et al., 2016). The models account for stellar rotation, parameterized according to the magnetic braking law of Matt et al. (2015); Amard et al. (2019). The initial rotation period following disc dissipation is set to 5.5 days, consistent with typical rotation rates observed in NGC 2362 (Irwin et al., 2008). The disc dissipation timescale is fixed at 6 Myr, based on Tu et al. (2015). Since the star-planet systems explored in this paper are older than 100 Myr, their present-day rotation rates are insensitive to the initial conditions; the same is true for rotation rates at the onset of IGW dissipation. We find that the critical magnetic field strength required for wave conversion, $B_{\mathrm{crit}}$, is sensitive to the grid resolution, which is determined by the parameter mesh_delta_coeff. Increasing the resolution (by decreasing mesh_delta_coeff) allows $B_{\mathrm{crit}}$ to converge at mesh_delta_coeff $\lesssim 0.5$, as we show in Fig.1. Therefore, we select mesh_delta_coeff =0.5=0.5 to balance computational efficiency and accuracy.
@@ 3.1 TOI-2458
Figure 3: Evolution of stellar rotation (top panel) and planetary orbital (bottom panel) periods for four star-planet systems with $P_{\mathrm{orb}}$ = 1, 1.5, 2.5, 2.5 days and $M_{\mathrm{pl}}=$ 10, 0.7, 2, and 4 $M_{\mathrm{J}}$, shown in red, orange, blue, and purple, respectively. The black dotted line depicts the rotation period evolution of a solitary star. Circles and crosses denote the onset of IGW breaking and Roche-lobe overflow (and planetary engulfment), respectively. The green square with error bars shows TOI-2458’s observed rotation period and age, along with their uncertainties from Šubjak et al. (2025).
Our hypothesis that IGW dissipation can explain the star’s anomalously fast rotation is supported by Fig. 3, illustrating the orbital evolution of four hypothetical HJs around TOI-2458, with planetary masses of 0.7, 2, 4, and 10 Jovian masses, and initial orbital periods of 1.5, 2.5, 2.5, and 1 day, respectively. These initial orbital periods represent the range of possible separations between the engulfed planet and its host star, considering the proximity of the hot Neptune TOI-2458 b. For the least massive HJs, such as those with masses as small as $0.3\>M_{J}$, the smaller initial separations above would place them below the Roche limit (for $M_{\mathrm{pl}}=0.3\>M_{\mathrm{J}}$ and $R_{\mathrm{pl}}=1.5\>R_{\mathrm{J}}$, equation (9) yields $a_{\mathrm{R}}\approx 0.026$ AU, which corresponds to an orbital period of $\approx 1.5$ days). Conversely, a planet with a larger initial separation would be too close to TOI-2458 b, as the mutual separation between the two planets would fall below 4 mutual Hill radii, which is the stability threshold for two-planet systems proposed by Wu et al. (2019). The mutual Hill radius $\Delta_{\mathrm{H}}$ in the units of mutual planetary separation is determined by:
@@
Rapid orbital migration, triggered by IGW breaking (the onset of which is shown by the black circles), modifies the stellar rotation period compared to that of a solitary star (shown with a black dotted line), as depicted in the top panel of Fig. 3. The bottom left panel reveals that the engulfment of planets with $M_{\mathrm{pl}}$ = 0.7 and 10 $M_{\mathrm{J}}$ occurs almost instantaneously after the onset of wave breaking. In contrast, 2.0 and 4.0 $M_{\mathrm{J}}$ planets undergo a prolonged phase of orbital migration that lasts more than 700 Myr.
A comparison between the migration of planets with $M_{\mathrm{pl}}$ = 0.7 and 4 $M_{\mathrm{J}}$ is particularly insightful. Despite the more massive planet depositing more angular momentum into stellar rotation, its orbital decay completes earlier, leaving more time for magnetic braking to erase the tidally enhanced rotation. Eventually, the rotation periods in both simulations converge, and at TOI-2458’s age, they reach values consistent with observations. The other two simulations shown in red and blue also allow the star to match the rotation period measurements. However, this time, the stellar rotation period equals its observed value at ages corresponding to either the lower limit from Šubjak et al. (2025) or an upper limit based on the TAMS age (which exceeds the upper age limit from Šubjak et al. (2025); given that other parameters such as $T_{\mathrm{eff}}$ and log g suggest that TOI-2458 is an MS star, here we use TAMS as the upper age limit).
@@ 3.2 GJ 504
TOI-2458 and GJ 504 represent two possible examples where efficient IGW dissipation leading to the HJ engulfment can explain the faster rotation rates of their host stars over expectations for single stars. They thus provide possible evidence for both the wave breaking scenario in a solar-type star (for TOI-2458) and for the magnetic wave conversion scenario in a hotter star with a convective core (GJ 504). Our hypothesis can be reinforced by exploring the presence of additional observational imprints of planetary engulfment. These signposts of planetary ingestion include, but are not limited to, metal enrichment of the stellar photosphere (e.g., Behmard et al., 2023; Liu et al., 2024; Soares et al., 2025; Lane et al., 2026, although see Soliman and Hopkins 2025; Sun et al. 2025 for alternative explanations), stellar spin misalignment relative to remaining planets (Tokuno, 2025), enhanced magnetic activity (Gehan, 2025), and increased luminosity (Montesinos et al., 2026).
@@
Finally, the results obtained in this section rely on the robustness of the adopted magnetic braking model from Matt et al. (2015). We note that the more recent magnetic wind parametrization by Spada and Lanzafame (2025) couples the braking torque to the moment of inertia of the convective envelope. Their parametrization suggests a weaker spin-down torque for stars in the mass range of TOI-2458. For stars with masses similar to GJ 504, however, both studies yield comparable estimates (see their Fig. 8). A reduction in magnetic wind braking for TOI-2458 would shift the black lines in Fig. 4 to the right, permitting convergence with the present-day rotation for more massive HJs at short orbital periods that finish their tidal decay earlier. At the same time, lower-mass planets would need to be initially located closer to the star, as their late engulfment would not leave the host enough time to spin down to the observed rotation rates (for instance, the 2 $M_{\mathrm{J}}$ planet shown in Fig. 3 will not be able to reproduce observations). Nevertheless, the uncertainties in the age of TOI-2458 exceed the differences between the wind models, precluding a detailed exploration of how this effect influences the size and shape of the parameter space region consistent with the present-day rotation.
@@ 4.5 Results
According to our synthetic populations, 12 – 13% of stars that once hosted a HJ and have not yet evolved off the MS have engulfed their HJ. The above fraction is calculated relative to all synthetic systems, including young and non-transiting ones. This finding is in good agreement with Lazovik (2023), although the latter study focused on solar-mass stars and considered a uniform SFH between 0 and 7 Gyr. The dependence of HJ engulfment probability on stellar mass is presented in Fig. 18. Furthermore, we can estimate the relative number of HJs producing transit timing variations detectable with modern facilities over 10 years of observations. The cumulative shift in transit times $T_{\mathrm{shift}}$ can be derived from the following (e.g. Wilkins et al., 2017):
@@
Figure 18: Engulfment probability (defined as the fraction of engulfed HJs prior to the TAMS or 10 Gyr) as a function of stellar mass. Red histogram: SFH from Mor et al. (2019). Black contour line: uniform SFH.
@@ 5.1 Can resonance locking explain observations?
In summary, our results suggest that the ‘pure’ application of the resonance locking mechanism — without considering non-linear effects — overestimates the migration rates of HJs, leading to an excess of short-period planets that is not seen in the data. The discrepancy between our conclusions and those of Millholland et al. (2025) likely stems from the limitations associated with fitting a universal scaling law for $Q^{\prime}$ in their study. We posit that this approach is likely insufficient to capture the complexity of the HJ population; instead, any robust empirical constraint on tidal theory must account for multiple migration regimes. Our findings imply that, while resonance locking may dominate during early stages or at larger separations, it must eventually transition into a wave-breaking regime to remain consistent with the observed orbital period distribution. This transition, particularly relevant during the later stages of the MS lifetime, underscores the necessity of testing future tidal models against both orbital period distributions and measured decay timescales simultaneously.
@@ 5.3 Other caveats
Another source of uncertainty is the contribution of planets that arrive at their present-day locations via high-eccentricity migration at late ages ($t>t_{\mathrm{gw}}$), which we refer to as ‘late-travelers’. Because our approach is designed to show a connection between the populations of young and old HJs, the late-stage implantation of planets would decrease the convergence between our synthetic and observed populations. Nevertheless, neglecting the contribution of these late-travelers is a reasonable simplification, as the dynamical instabilities that drive high-eccentricity migration are expected to take place relatively early in a system’s evolution. This is supported by observations of hot Neptunes, which were found to be more abundant in systems with ages between 100 Myr and 1 Gyr (Fernandes et al., 2025). It is possible, though, that a fraction of HJs are formed through secular interactions between planets on longer timescales, as discussed in e.g. Hamer and Schlaufman (2022); Schmidt and Schlaufman (2026).
@@
While the ages for the majority of our systems were computed using isochrone fitting based on the homogeneous grid by Swastik et al. (2024), the subsample taken from the NASA Exoplanet Archive includes an admixture of different isochrone models and age-determination techniques. This diversity could potentially introduce systematic biases into our analysis. Nevertheless, typical age uncertainties in our sample likely exceed these possible biases, which is why we expect their impact on our overall conclusions to remain small. Future studies would benefit from repeating this procedure as larger samples of HJ systems with homogeneously determined ages become available.
@@ 6 Conclusions
  • Up to 13% of stars that once hosted a HJ and have not yet evolved off the MS have engulfed their planet.
@@ Acknowledgments
We gratefully acknowledge Dr. Chowbay Swastik for the provided data on ages and other parameters of star-planet systems. YL thanks Anastasiya Yarovova for helpful discussions. We would also like the thank the referee for their careful reading and constructive suggestions. This research has used the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.